OpenStudy (anonymous):

x squared -1 over 1-x , I'm suppose to simplify but I dont see how, please help me understand

OpenStudy (anonymous):

\[x^2-1/1-x\] I'm going to assume this is the equation you are asking about. So what you do is you factor the top polynomial. \[x^2-1=(x+1)(x-1)\] Thus, you have (x+1)(x−1)/(1-x)

OpenStudy (anonymous):

\[\LARGE x^2-\frac{1}{1-x}\] this ? or like post above :)

OpenStudy (anonymous):

But you can rewrite x-1 as -(1-x) so you have...

OpenStudy (anonymous):

i did the first one you posted

OpenStudy (anonymous):

-(x+1)(1-x)/1−x=-(x+1) so your final answer is -x-1

OpenStudy (anonymous):

 is what i did

OpenStudy (anonymous):

Im sorry i stil dont see how you got -x-1 from that

OpenStudy (anonymous):

Okay so \[(x^2-1)/(1-x)=(x+1)(x-1)/(1-x)=-(x+1)(1-x)/(1-x)\]

OpenStudy (anonymous):

You cancel out (1-x) then you are left with -(x+1)=-x-1